System Status

1System Status

Reported Date / Time 16/09/2016 09:40
Status Active
Issue Notice to All Users

All future System Status Page Updates will be displayed on the RFCA Home Page
Reported Date / Time 03/09/2016 10:40
Status Resolved
Resolved Date / Time 07/09/2016 11:00
Issue All PH2 Users

The in-coming and out-going E-Mails from with PH2 (RFCA.MOD.UK) appear to be working correctly now; therefore this System Status has been resolved - however, SPOC will continue to monitor the situation.

05/09/2016 08:40 - The issue concerning E-Mails being received and sent from the MOD appears to be resolved - and E-Mail are moving in and out at the normal pace. SPOC will continue to monitor.

There is currently an issue with E-Mail traffic being received and sent into and out of the MOD System. SPOC are currently unable to resolve this issue as it is in the hands of CASC (MOD): however, SPOC will monitor the situation.
Reported Date / Time 02/09/2016 14:00
Status Resolved
Resolved Date / Time 07/09/2016 11:00
Issue iPhone Security Issue

All Users of issued iPhones are to carry out the Security Update iOS 9.3.5 straight away without fail. Apple have identified a major security issue that this patch will fix.
Reported Date / Time 30/08/2016 07:50
Status Resolved
Resolved Date / Time 30/08/2016 09:27
Issue Mail from PH2e to PH2 being blocked.

Mail from PH2e to PH2 is being blocked with an error (554 Blocked).
SPOC is currently investigating and will resolve this as soon as possible.
Reported Date / Time 17/08/2016 09:40
Status Resolved
Resolved Date / Time 26/08/2016 15:55
Issue All PH2 Users

User may see a message stating Group Policy Update Log Off, please select no and carry on as normal.
Reported Date / Time 08/08/2016 08:20
Status Resolved
Resolved Date / Time 09/08/2016 08:00
Issue PH2 SharePoint Issues

08/08/2016 08:55 - There appears to be the same issues with SharePoint that are affecting EMSP; other RFCA's may be experiencing the same issues; if so please report them to SPOC.

There are currently 2 issues concerning South East RFCA SharePoint and Greater London SharePoint Sites; SPOC are currently investigating these issues and will update the System Status Page when we have further information.
Reported Date / Time 03/08/2016 08:30
Status Resolved
Resolved Date / Time 03/08/2016 08:40
Issue PH2 New

There appears to be an issue for some users not being able to open the home page for PH2 New. If you experience this please notify your RSA or email SPOC ( if you have access to email on PH2e or someone near you does.

SPOC are investigating this issue.
Reported Date / Time 29/07/2016 08:10
Status Resolved
Resolved Date / Time 29/07/2016 13:15
Issue PH2 FileStore 1

29/07/2016 08:35 - LO / HI / WM should be able to access their data now, please log off and back on if unable to do so.

There is currently an issue with PH2 Filestore1 which is being investigated by SPOC. Users from EM / NE / NW / WA / WM / WX / YH will not have access to their N drive archive data for the time being.

SPOC is working to resolving this issue as soon as possible; further updates will be posted here.
Reported Date / Time 28 July 2016 12:45
Status Resolved
Resolved Date / Time 29/07/2016 07:50
Issue PH2 Users

The migration of data from the old PH2 system to the new will take place today between 1300 and 1330 hours. If any user experiences problems/issues with their Outlook during this time, please close Outlook and wait until after 1330 hours before attempting to connect Outlook again.
Reported Date / Time 21/07/2016 09:00
Status Resolved
Resolved Date / Time 22/07/2016 07:35
Issue All PH2e and PH2 Users

Due to the on-going issues with BT Internet, the SPOC telephone lines have been affected; therefore if you require assistance please ring 01252 357619 until further notice.
Reported Date / Time 20/07/2016 12:20
Status Resolved
Resolved Date / Time 20/07/2016 13:00
Issue Symphony Users

We need to restart SYMPHONY Server to fix an issue with the email component, can everyone please log out between 12.30 and 13.00. In that period you can still access Symphony from Web Apps.

RSAs please cascade this message as required.
Reported Date / Time 20/07/2016 10:15
Status Resolved
Resolved Date / Time 29/07/2016 08:00
Issue Internet Access on PH2e and PH2

22/07/2016 07:35 - BT have managed to address and resolve some of their issues concerning the major outage they experienced a couple of days ago: however, some issues are still outstanding and some PH2 users may still have issues connecting to some Internet site.

There is currently a wide BT Issue that is affecting Internet Access to some sites. SPOC is currently investigating and in talks with BT; when further updates are available they will be displayed here.
Reported Date / Time 19/07/2016 13:50
Status Resolved
Issue PH2 New

There was a server error within PH2 New; this has been investigated and resolved by SPOC; all personnel should now be able to log back into PH2 New.
Reported Date / Time 14/07/2016 12:25
Status Resolved
Resolved Date / Time 14/07/2016 12:55
Issue Notice to Lync Users

The Lync Server is now working correctly, and user should be able to use Lync.

There is currently a certificate error with the Lync Server; SPOC are investigating. Once resolved, the System Status Page will be updated.
Reported Date / Time 13/07/2016 07:55
Status Resolved
Resolved Date / Time 15/07/2016 09:10
Issue East Midland & Yorkshire and The Humber RFCA's

All East Midlands, York & Humber RFCAs users must follow instruction for PH2 email migration as detailed in the User Migration instruction para 3. Please see link below:

PH2 Migration User Instructions
Reported Date / Time 11/07/2016 11:25
Status Resolved
Resolved Date / Time 15/07/2016 09:10
Issue Notice to All User:

SSL Renewal – 14/07/2016 - Please note the Security Certificate on PH2 London site will be renewed on Thursday evening and there will be an outage of a few minutes during which users will not be able to login. This does not affect the users that have been migrated to PH2 New
Reported Date / Time 04/07/2016 14:05
Status Resolved
Resolved Date / Time 05/07/2016 07:55
Issue PH2 Users

All PH2 Users are currently experiencing an issue with Outlook; SPOC are aware and working on a solution.
Reported Date / Time 30/06/2016 08:50
Status Resolved
Resolved Date / Time 04/07/2016 07:50
Issue PH2 Users

30/06/2016 11:02 - Some PH2 Users may still experience issues with their profile. Can you please log out of the vdt. In the interim can you please use the Web Bookmarks.

30/06/2016 10:00 - can all PH2 Old and PH2 New Users log out of the PH2 system, if not already done so, and then login again. You should be given your correct profile now.

30/06/2016 09:10 - Can all PH2 New users log out of the system and then sign back in again; this should allow you to access the New PH2 system with your correct profile.

30/06/2016 09:05 - Until SPOC and resolved the Temporary Profile Issues, all users are advised to use the Web Bookmarks on PH2 Old and Web Application on PH2 New.

Currently the PH2 System is giving people Temporary Profiles; SPOC is currently working on the issue - and will update this Status Page with further updates.
Reported Date / Time 29/06/2016 09:50
Status Resolved
Resolved Date / Time 04/07/2016 07:50
Issue PH2 New

29/06/2016 10:15

Please use the Web Applications for SharePoint and Symphony until SPOC have resolved the PH2 New performance issue.

Currently the PH2 New system is experiencing performance issues; SPOC 3rd Line Technicians and SPOC Network Team in connection with relevant resources are investigating; further updates will be posted onto the System Status Page once known.
Reported Date / Time 21/06/2016 11:15
Status Resolved
Resolved Date / Time 04/07/2016 07:50
Issue Westminster Users

The Westminster Team have been informed of the issue, where users gets a message ‘Client is not trusted’ - they are working on it and this system status entry will be updated once we have confirmation that the issue has been resolved’

Please note RFCA IT Single Point of Contact (SPOC) is available on 01252 413719 at the following times

– except public holidays, for example Christmas day

0800-1800: Mon-Fri (Full support services)
1800-2200: Mon-Fri (Telephone support only)
1000-1600: Weekends (Telephone support only)
Or by email at any time.